1. A container has 500.0 mL of an aqueous solution containing OH– ions. How many grams of OH– ions are in the solution if the molarity is 0.696 M?

2. Cylinder 1 contains a 1.50 M stock solution of Cu(NO3)2. What volume of stock is needed to prepare 20.0 mL of 0.0750 M Cu(NO3)2 solution shown in Cylinder 2?

I am so confused! please help
Answer #1

Answer (1) We know that

Molarity= moles of solute / volume of solutioin in L

Here Molarity= 0.696 M and Volume of solution = 500 mL = 0.5 L

rearranging the formula and gettig moles of solute

moles of solutre = Molarity* Volume of solution = 0.696*0.5= 0.348 moles of OH- ions

mass of OH- = moles*molar mass = 0.348*17.008 g/mol= 5.918 grams OH-